### v2

# Empirical Analysis of Transaction Conflicts in Ethereum and Solana for Parallel Execution

Abstract
This paper presents a comprehensive analysis of historical data across two popular blockchain networks: Ethereum and Solana. Our study focuses on two key aspects: transaction conflicts and the maximum theoretical parallelism within historical blocks. We aim to quantify the degree of transaction parallelism and assess how effectively it can be exploited by systematically examining block-level characteristics, both within individual blocks and across different historical periods. In particular, this study is the first of its kind to leverage historical transactional workloads to evaluate transactional conflict patterns. By offering a structured approach to analyzing these conflicts, our research provides valuable insights and an empirical basis for developing more efficient parallel execution techniques for smart contracts in the Ethereum and Solana virtual machines. Our empirical analysis reveals that historical Ethereum blocks frequently achieve high independence, over 50% in more than 50% of blocks, while Solana historical blocks contain longer conflict chains, comprising $\sim$59of the block size compared to $\sim$18in Ethereum, reflecting fundamentally different parallel execution dynamics.
1 Introduction
Blockchain virtual machines (VMs) execute a block of user-defined transactions, each performing a sequence of read and write operations on the states of user accounts. A block proposer, a validator node in the network, takes as input a block consisting of $n$ transactions and a preset total order $T_{1}\rightarrow T_{2}\ldots,\rightarrow T_{n}$ that enforces the state of transactional execution within that block. The node attempts to execute, in parallel, the $n$ transactions such that the state resulting from the parallel execution of the $n$ transactions must be the state resulting from the sequential execution of transactions $T_{1}\cdot T_{2}\cdots T_{n}$. Ensuring consistency across the network requires that all nodes execute transactions in the preset order, referred to as the preset serialization [7, 11].
VMs utilize different execution strategies that directly affect transaction throughput. Ethereum [4] and Solana [16] represent two distinct architectures, each employing a different transaction execution approach. Ethereum processes transactions sequentially without prior knowledge of the states that they access. The state accessed by a transaction is determined only at the time of block execution. We refer to this as a read-write oblivious execution model, where read-write sets of the account states the transaction accesses are not known a priori. In contrast, Solana enables parallel execution by requiring clients to specify state access information for each transaction, which helps to improve throughput by supporting parallel execution of the smart contract transactions. We refer to this as the read-write aware execution model.
When block transactions are executed, they may modify the same account address or storage slot in a smart contract, potentially resulting in conflicts if one transaction accesses data modified by another. A conflict occurs when two or more transactions access the same state and at least one of them performs an update operation. The transactions in the block must be executed sequentially in preset order, the order of the transactions in the block, or they must be serializable to the preset order if executed in parallel, to provide deterministic execution across the distributed network. The order in which conflicting transactions are executed directly affects the final state. Two transactions are independent if they do not modify the same state, and the outcome of one does not impact the other. The independent transactions can be executed in any order or in parallel, as their execution does not interfere with other transactions.
The longest chain of conflicting transactions is one of the parameters that determines the maximum theoretical limit (or minimum execution time) for parallelization within a given block, including the number of conflicting transactions, number of conflicts and different conflict sets (conflict families). A conflict family is a group of transactions that are mutually dependent on shared states. Identifying conflict families is essential for optimizing transaction execution, as it helps in detecting hotspots and employing the best suitable parallel execution strategy.
Potential transaction conflicts can be identified through static analysis of smart contract bytecode, while exact conflicts can be detected via optimistic execution by analyzing the read and write sets of each transaction. Efficient conflict identification and resolution with minimal overhead, including abort and re-execution costs, are crucial for enhancing performance. The timing and cost of conflict detection and resolution significantly affect the efficiency of parallel execution. The distribution of conflicts and state access patterns within historical blocks is the main focus of this study, along with the metrics we discussed above that directly or indirectly affect the parallel execution of transactions and overall throughput. These insights support the development of techniques that improve execution efficiency and enhance parallelization across blockchain networks.
Previous studies [10, 13] analyzing historical Ethereum blocks have demonstrated promising potential for parallel execution.
Paper Contributions: In this paper, we analyze two blockchain networks with different transaction execution models to better understand their conflict distributions and impact on performance. We present a detailed study of conflict relationships in these two most popular smart contract execution paradigms. We also present a comparative study of transactional conflict patterns in Ethereum versus Solana to empirically assess how their read-write oblivious and read-write aware execution models influence the potential for parallel execution. Our findings offer insights into the ground truth for the maximum parallelism that can be extracted in historical transactions, constrained by inherent conflicts within the executed blocks.
Our analysis of Ethereum blocks reveals that more than 50% of the blocks consistently contain over 50% independent transactions, with over 94% exceeding the 40% threshold in recent periods, indicating a growing opportunity for parallel execution. In contrast, recent historical blocks in Solana exhibit significantly longer conflict chains ($\sim$59of block size vs. Ethereum’s $\sim$18%), fewer conflict families (28 vs. 72) and only 4% independent transactions compared to Ethereum’s 51%, highlighting the dense dependency patterns in Solana’s historical workloads.
Paper Outline: In Section 2, we present the background and motivation of our conflict study. The system model and formal conflict definitions are provided in Section 3. Section 4 introduces the conflict metrics used to evaluate parallelism in historical blocks and describes the data extraction process. The empirical analysis of Ethereum and Solana are presented in Section 5 and Section 6, respectively. A comparative study of the conflict patterns between Ethereum and Solana is provided in Section 7. Finally, we conclude in Section 8 and provide some expanded details of our analysis in the Appendix.
2 Motivation
As blockchain systems evolve to support higher throughput, parallel execution of transactions becomes critical for scalability. However, the effectiveness of parallel execution is fundamentally constrained by transaction conflicts where two or more transactions access the overlapping state, especially when at least one access is a write. Ethereum [4] and Solana [16] represent two contrasting execution paradigms. Ethereum follows a single-threaded, sequential execution model, which we call the read-write oblivious execution model. While Solana employs parallel execution based on explicit read-write set declarations, called the read-write aware execution model, enabling parallelization based on declared state access.
In the read-write aware model, the read-write sets of transactions are known prior to execution. This enables the static construction of a conflict matrix by analyzing the intersections of read and write sets, allowing for informed scheduling and conflict-free parallel execution. This enables the formation of a conflict graph that captures dependencies between transactions, allowing for optimal or near-optimal scheduling of non-conflicting transactions across multiple processing threads. Independent transactions, those with disjoint read and write sets, can be safely executed in parallel, significantly improving throughput. The conflict families and longest conflict chains identified through this matrix further guide the execution engine in minimizing serialization. This static approach results in deterministic execution with minimal overhead as rollback mechanisms are rarely needed. Consequently, systems that can statically derive read-write sets (e.g., via access lists or symbolic analysis) could be well-positioned to exploit parallelism in blockchain environments. However, accurately determining these access sets can introduce additional computational overhead, especially for contracts with complex or data-dependent behaviors. Including explicit access lists within transactions increases the transaction size, thus increasing bandwidth and storage requirements across the network.
In the read-write oblivious model, when read-write sets are not known until runtime, the execution engine must dynamically detect conflicts during transaction processing. In this case, the conflict matrix is incrementally built based on observed state accesses, often necessitating speculative execution strategies such as software transactional memory (STM) or optimistic concurrency control. Transactions are executed in parallel under the assumption of independence, but if conflicts are detected, such as overlapping writes or read-write dependencies, transactions must be rolled back and re-executed. This approach introduces execution overhead, increases nondeterminism, and limits achievable parallelism, especially in workloads with high conflict density. Furthermore, to preserve determinism and ensure that every node in the network maintains a consistent global state, parallel execution must adhere to a predefined serialization order, denoted as preset serialization, which serves as the canonical transaction sequence. This requirement further constrains execution flexibility and highlights the trade-offs between generality and performance in conflict-agnostic parallel execution models.
[…]
Figure 1: Execution in Ethereum (read-write oblivious) versus Solana (read-write aware).
For example, Figure 1 illustrates Ethereum’s read-write oblivious execution and Solana’s read-write aware execution using a block $B_{i}$ containing three transfer transactions. In the existing Ethereum model (see Figure 1a), transactions $T_{1},\ T_{2},\ \text{and},\ T_{3}$ are executed sequentially in the preset order. Any parallel execution of these transactions in the EVM requires optimistic execution because the read and write sets are unknown prior to execution, making conflict detection possible only at runtime. Consequently, if conflicting transactions are executed speculatively in parallel, they must be validated and potentially rolled back to satisfy the preset serialization, adding significant execution overhead. For example, if $T_{3}$ executes before $T_{2}$, it must be aborted and re-executed after $T_{2}$ commits (tryC2). In contrast, in the Solana model transaction’s read and write sets are known in advance before execution, as shown in Figure 1b. This allows the runtime to detect that $T_{2}$ and $T_{3}$ conflict on $X_{3}$ and $X_{4}$, while $T_{1}$ is independent. As a result, read-write aware execution can safely execute $T_{1}$ in parallel with $T_{2}$ and defer $T_{3}$, demonstrating how this model allows for efficient and conflict-free parallel execution.
Research Goals: In this paper, we seek to understand the limits of parallelism in Ethereum (read-write oblivious) and Solana (read-write aware) by conducting a comprehensive empirical study of the nature of transactional conflict patterns. Several studies have analyzed the conflict structures in historical Ethereum blocks to understand conflicts and assess the potential for parallelism [6, 9, 10, 13, 20]. These works typically focus on static conflicts, access patterns, optimistic concurrency control, or speculative execution. However, they do not compare the observed conflict behavior with chains such as Solana, where the virtual machine executor is given input as a batch of transactions with their read-write sets.
To our knowledge, no previous study has performed a comparative historical analysis of transaction conflicts on Ethereum [4] and Solana [16], representing read-write oblivious and read-write aware execution models, respectively. This leaves a critical gap in understanding how execution design choices shape conflict characteristics and thus the realizable throughput from parallelism. Such an analysis allows us to present answers to the following research questions:
RQ1
Quantify what fraction of transactions are independent and how transactions distributed across conflict families?
RQ2
What are the typical depth and width characteristics of conflict chains in historical blocks?
RQ3
How does the maximum achievable parallelism compare between read-write aware execution (as in Solana) and read-write oblivious execution (as in Ethereum)?
RQ4
What are the fundamental design trade-offs in block transactional execution across different virtual machines, particularly the EVM and the SVM?
By analyzing historical block traces from both chains, we aim to identify conflict families, independent transactions, and structural patterns across blocks. This will help to design conflict-aware scheduling, adaptive execution models, and hybrid virtual machine architectures that achieve higher throughput under realistic workloads. It also provides empirical insight into the frequency, structure, and granularity of conflicts, enabling us to assess the potential gains and limitations of parallelism.
3 System Model
This section formalizes the execution semantics of the blockchain virtual machine (BVM) from a block-execution-centric perspective, defining conflicts between transactions in a block. The BVM definition provides a foundation, while the EVM and Solana VM (SVM) formalizations highlight differences in state management and conflict definitions.
Blockchain Transaction Execution: A blockchain virtual machine (BVM) is a deterministic state machine that processes transactions and updates the global state according to predefined transition functions.
Formally, a BVM is defined as a tuple:
$\mathcal{BVM}=(\mathcal{S},\mathcal{B},\mathcal{E},\mathcal{R})$ (1)
where:
  • $\mathcal{S}$ is the set of all possible states of the blockchain.
  • $\mathcal{B}$ is the set of all possible blocks, each containing a list of transactions $\mathcal{T}$.
  • $\mathcal{E}:\mathcal{S}\times\mathcal{B}\to\mathcal{S}$ is the execution function that maps a state and a block to a new state.
  • $\mathcal{R}:\mathcal{S}\times\mathcal{B}\to\{0,1\}$ is the validity function that determines whether a block is valid in a given state.
[…]
Transactional Conflicts: Conflicts in a BVM arise when two transactions $T_{i},\ T_{j}\in B$ (where $B$ is a block) access a shared state $s\in\mathcal{S}$ and at least one of them modifies it. This condition implies that the execution order of the transactions affects the final state if the transactions are executed in parallel or reordered within the block $B$.
Formally, a conflict occurs if:
$\exists s\in\mathcal{S},B=\{T_{i},T_{j},\dots\},\quad\mathcal{E}(\mathcal{E}(s,T_{i}),T_{j})\neq\mathcal{E}(\mathcal{E}(s,T_{j}),T_{i})$ (2)
Indicating that the execution of $T_{i}$ and $T_{j}$ is order-dependent within a block and are conflicting. Equivalently, we say that two transactions $T_{i}$ and $T_{j}$ conflict if there exists a common VM state accessed by both and it is contained in write-seti or write-setj or both. [^1] We remark that, in practice, it might be not be possible to distinguish between read and write accesses performed within transactions a priori. This is the case, for instance, in the read-write oblivious model of Ethereum. A historical analysis of blockchain data enables empirical evidence of conflict percentages and the extent of parallelism that can be exploited in real-world transaction interactions.
Ethereum Virtual Machine (EVM) processes transactions in a shared global state. Formally, it can be defined as:
$\mathcal{EVM}=(\mathcal{S}_{E},\mathcal{B}_{E},\mathcal{E}_{E},\mathcal{R}_{E})$ (3)
where:
  • $\mathcal{S}_{E}$ is the state that comprises account balances, contract storage, and program code.
  • $\mathcal{B}_{E}$ is the set of blocks, each containing a list of transactions $\mathcal{T}_{E}$.
  • $\mathcal{E}_{E}:\mathcal{S}_{E}\times\mathcal{B}_{E}\to\mathcal{S}_{E}$ is the execution function, following a stack-based computation model.
  • $\mathcal{R}_{E}$ ensures block validity based on nonce ordering, gas availability, and state transitions.
Conflicts in the EVM occur because of shared state dependencies within a block. Two transactions $T_{i},T_{j}\in B_{E}$ conflict if they access the same account or contract storage location and at least one of them update it. Formally, a conflict exists if:
[…]
$\quad\textit{write}_{i}(x)\ \vee\ \textit{write}_{j}(f)\ \wedge\ x\in\textit{Dset}(T_{j})\vee f\in\textit{Dset}(T_{j})$
where $X$ is the set of accounts and $\mathcal{F}$ is the set of storage fields.
This can be decomposed into two cases:
  • EOA Conflict: Two transactions $T_{i}$ and $T_{j}$ conflict if they both access a common externally owned account (EOA) address and at least one modifies its balance or nonce.
  • Smart Contract Storage Conflict: Two transactions $T_{i}$ and $T_{j}$ conflict if they access a common storage location within a contract address and at least one of them writes to it.
Solana Virtual Machine (SVM) requires each transaction explicitly declares its read and write sets (unlike the EVM), allowing the system to schedule non-conflicting transactions concurrently through it’s SeaLevel [21] execution engine. Formally, SVM defined as:
$\mathcal{SVM}=(\mathcal{S}_{S},\mathcal{B}_{S},\mathcal{E}_{S},\mathcal{R}_{S})$ (5)
where:
  • $\mathcal{S}_{S}$ is the global state, consisting of account data and contract state.
  • $\mathcal{B}_{S}$ is the set of blocks, each containing a list of transactions $\mathcal{T}_{S}$, with associated read sets (Rset) and write sets (Wset).
  • $\mathcal{E}_{S}:\mathcal{S}_{S}\times\mathcal{B}_{S}\to\mathcal{S}_{S}$ is the parallel execution function. Given the global state and a block, it applies a schedule over the transactions $\mathcal{T}_{S}$ using access-set profiling to determine a non-conflicting execution order. Internally, this schedule respects conflicts defined by overlaps in Rset and Wset.
  • $\mathcal{R}_{S}$ ensures block validity, including signature checks and state consistency.
Conflicts in the SVM arise when transactions within the same block attempt to modify overlapping data. Formally, two transactions $T_{i},\ T_{j}\in B_{S}$ conflict if:
$(\textit{Wset}{{}_{i}}\cap\textit{Wset}{{}_{j}})\cup(\textit{Wset}{{}_{i}}\cap\textit{Rset}{{}_{j}})\cup(\textit{Rset}{{}_{i}}\cap\textit{Wset}{{}_{j}})\neq\emptyset$ (6)
This ensures that transactions in a block are only executed concurrently if they do not interfere with each other’s read and write sets.
4 Defining and Extracting Conflicts Metrics
In this section, we first introduce the conflict metrics used in our analysis to evaluate parallelism in historical blocks. We then discuss how we extracted the data for the conflict analysis.
Let $T_{i}$ denote a transaction. Each transaction is associated with:
  • A read-seti (Rseti) — state keys read by $T_{i}$.
  • A write-seti (Wseti) — state keys written by $T_{i}$.
In the read-write aware model, each transaction explicitly specifies its read-write sets prior to execution. In the read-write oblivious model, read-write sets are computed dynamically at runtime through speculative execution to detect potential conflicts and enable parallel execution.
We illustrate the conflict metrics using the following set of transactions {$T_{1},T_{2},\dots,T_{8}$}. For simplicity, we consider small read-write sets, although in practice these sets can be significantly larger depending on the blockchain network.
[…]
1. Conflict dependent transactions:
Two transactions $T_{i}$ and $T_{j}$ are said to be conflict dependent if:
  • $T_{i}$ precedes $T_{j}$ in the preset order.
  • $T_{i}$ and $T_{j}$ in conflict:
    $\text{Conflict}(T_{i},T_{j},s)=\begin{cases}1&\text{if }s\in(\textit{Wset}{{}_{i}}\cap\textit{Wset}{{}_{j}})\cup(\textit{Wset}{{}_{i}}\cap\textit{Rset}{{}_{j}})\cup(\textit{Rset}{{}_{i}}\cap\textit{Wset}{{}_{j}})\\ 0&\text{otherwise}\end{cases}$ (7)
    It means that the read-write sets of $T_{i}$ and $T_{j}$ have a non-empty intersection on at least one conflicting state access.
    Example: $T_{2}$ and $T_{3}$ are in write-read and write-write conflict because Wset2 = {$X_{3}$, $X_{4}$} and Rset3 = {$X_{3}$}, as well as Wset3 = {$X_{3}$, $X_{6}$}, hence overlap, as a result $T_{3}$ conflict dependent on $T_{2}$. Similarity $T_{3}-T_{4}$ and $T_{5}-T_{6}$ are conflict dependent.
2. Conflict independent transactions:
The number of independent transactions in the block measures how much of the workload can be executed in parallel. A higher percentage of independent transactions suggests that the blockchain can scale better with more cores. Let B = $\{T_{1},T_{2},\ldots,T_{n}\}$ be the set of transactions in a block. Let $I\subseteq B$ be the set of independent transactions such that:
$\forall T_{i}\in I,\forall T_{j}\in B,i\neq j:\quad T_{i}\text{ and }T_{j}\text{ do not conflict}$
Then the percentage of independent transaction is defined as: $\frac{|I|}{|B|}\times 100\%$.
Example: $T_{1}$, $T_{7}$, and $T_{8}$ are conflict-independent transactions, since read-read is not a conflicting operation and there is no overlap with other writing transactions. The independent percentage in our example set of transactions is $\frac{3}{8}\times 100\%=37.5\%$.
3. Longest chain of conflicting transactions:
The longest conflict chain, or the longest chain of dependent transactions, represents the longest sequence of transactions where each transaction conflicts with the transaction next in the preset order. It quantifies minimal execution time under parallelism and bounds the theoretical maximum speed-up based on the length of sequential dependencies. Even with the number of available resources (threads/cores), the algorithm must execute the transactions in the chain sequentially because of their interdependencies. The longer the chain, the lower the maximum achievable speedup.
To compute the longest chain of conflict-dependent transactions, construct a conflict graph $G=(V,E)$, where each vertex $v_{i}\in V$ represents a transaction $T_{i}$, and there is an edge $(v_{i},v_{j})\in E$ if $T_{i}$ and $T_{j}$ conflict. Then, the longest chain of conflicting transactions is:
Longest chain of $\text{Longest chain of conflicts}=\max_{P\in\text{Paths}(G)}|P|$
Example: There are two conflict chains: $c_{1}$: $T_{2}$to $T_{3}$to $T_{4}$ and $c_{2}$: $T_{5}$to $T_{6}$. The longest chain is $c_{1}$, since $|c_{1}|$ $>$ $|c_{2}|$.
4. Conflict families:
The conflict families cluster together transactions that conflict with each other and helps in identifying serializable groups. It helps in measuring parallelism where independent families can be executed in parallel, whereas transactions within the same family must be serialized due to potential conflicts.
A conflict family can be viewed as a connected component in the conflict graph $G$. That is, a set of transactions $F\subseteq B$ such that:
  • $\forall T_{i},\ T_{j}\in F,\ \exists\text{ a path in }G\text{ between }v_{i}\text{ and }v_{j}.$
  • $F$ is maximal, i.e., no larger set including $F$ satisfies the above.
Formally, $F_{1},F_{2},F_{3},\dots$ are all connected components of $G$, i.e., $\{F_{1},F_{2},\ldots,F_{k}\}$ denote such conflict families that are disjoint.
$B=\bigcup_{i=1}^{k}F_{i},\quad F_{i}\cap F_{j}=\emptyset\text{ for }i\neq j$
Example: The following are the conflict families in the transaction set of our example:
  • $F_{1}$: $\{T_{1}\}$
  • $F_{2}$: $\{T_{2},T_{3},T_{4}\}$
  • $F_{3}$: $\{T_{5},T_{6}\}$
  • $F_{4}$: $\{T_{7}\}$
  • $F_{5}$: $\{T_{8}\}$
There are five conflict families; although both $T_{7}$ and $T_{8}$ read the same state $X_{11}$, they do not conflict, as read-read operations are non-conflicting, and therefore belong to separate conflict families.
5. Most dense conflict family:
The most dense conflict family is the one with the most transactions: $F^{*}=\arg\max_{F_{i}}|F_{i}|$. It sets a theoretical upper bound on execution speedup, meaning, the size and density of the largest conflict family limit the achievable parallelism:
  • More dense $\rightarrow$ more serialization $\rightarrow$ less parallelism.
  • Less dense $\rightarrow$ more independent execution $\rightarrow$ higher parallelism.
In our example, the most dense conflict family is $F_{2}$ with 3 transactions.
6. Total and Write-write conflicts:
The total conflicts measures how many transaction pairs access overlapping state in conflicting ways (read-write or write-write), indicating the general level of transaction interdependency. Higher total conflicts mean higher contention, suggesting lower parallelism without advanced conflict management. On the other hand, write-write conflicts are stronger conflicts, two transactions trying to modify the same state, and almost always require serialization. However, multi-versioning allows transactions to proceed in parallel by creating separate versions of conflicting writes, thereby reducing blocking and increasing parallelism even in the presence of write-write conflicts. In our analysis, total conflicts are determined by individual state accesses among transactions in preset order where each pair of conflicting accesses contributes one conflict.
$\text{Total conflicts}=\sum_{1\leq i<j\leq n}\ \sum_{s\in\mathcal{S}}\mathbf{1}\ (s\in\textit{Rset}{{}_{i}}\cap\textit{Wset}{{}_{j}}\vee s\in\textit{Wset}{{}_{i}}\cap\textit{Rset}{{}_{j}}\vee s\in\textit{Wset}{{}_{i}}\cap\textit{Wset}{{}_{j}})$
The write-write conflicts are the conflicts at the level of individual states written by the transactions.
[…]
Where: $\mathcal{S}$ is the set of states (smart contract storage slots, accounts, etc.), $n$ is total number of transactions in the block.
Example: There are six total conflicts in our example and three write-write conflicts as follows:
  • $T_{2}$
    $\to$ $T_{3}$: write-read and write-write conflict on state $X_{3}$.
  • $T_{3}$
    $\to$ $T_{4}$: write-read and write-write conflict on state $X_{6}$.
  • $T_{5}$
    $\to$ $T_{6}$: write-read and write-write conflict on state $X_{9}$.
For these conflict metrics, we analyze transactions from different historical periods (HPs)—defined as contiguous time windows on the blockchain timeline (e.g., grouped by block ranges or timestamps) during which blocks are analyzed. Different historical periods exhibit varying characteristics in terms of transaction load and state access patterns. Therefore, to capture the temporal dynamics patterns of transaction conflict, we performed our analysis over multiple historical periods on both blockchain networks.
Extracting Block Conflict Data
Ethereum
In our analysis, we segregate Ethereum transactions into two types: ETH transfer transactions and smart contract transactions, to analyze and understand the access and conflict patterns. ETH transfer transactions perform pure value transfers between externally owned addresses (EOAs) or to smart contract addresses. On the other hand, smart contract transactions interact with sender addresses and contract address(es) to modify blockchain state via function calls and storage updates within the contract(s).
We trace the states accessed by transactions within a block by using the callTracer and prestateTracer [2], which provides a full view of the block’s pre-state, the state required for the execution of current block. We identify all EOAs, contract addresses, and storage locations within contracts that are accessed by examining transaction data and the pre-state for it. This enables the detection of potential conflicts arising from overlapping state modifications by block transactions. As defined earlier in Equation 4, two EVM transactions $T_{i}$ and $T_{j}$ are in conflict if: (1) $T_{i}$ and $T_{j}$ both access a common EOA address; or (2) $T_{i}$ and $T_{j}$ both access a common storage location within the contract address(s).
It should be noted that two transactions are considered independent (non-conflicting) if they are initiated by separate EOAs and access different addresses and storage locations within all the contracts that they access. However, there is a special case in which every transaction updates the Coinbase account (block proposers account) for fee payment. As a result, all transactions are logically in conflict, unless the Coinbase account is treated as an exception. For this reason, when we analyze conflicts, we remove the Coinbase account from the transactions access-set (or Dset). In [12], a solution is proposed to collect the fee payment for each transaction independently, and updates to the Coinbase account are cautiously at the end of the block, allowing transactions to be executed in parallel.
We use an exclusive-access paradigm for our analysis of transaction conflict in Ethereum blocks. Since read-set information is not explicitly provided with prestateTracer [2] and transaction accessList [1], we treat each operation as an exclusive update operation, which may result in overestimating conflicts. For example, if $T_{i}$ and $T_{j}$ access the same state, they will conflict in the current analysis; however, there could be no conflict if they both just read the state in practice. Therefore, a more detailed analysis that separates accesses into read and write operations would likely reduce the conflict numbers presented in the Section 5.
Solana
In contrast to Ethereum, Solana transactions are made up of the account-access specification, a list of accounts to read (read-set) or write (write-set[16]. This specification is added to the transactions upfront by the clients through RPC node interaction. The success or failure of a transaction depends on the freshness of its read-write sets, from the moment it is added to the transaction by the client until its execution at the validator node. Two Solana transactions $T_{i}$ and $T_{j}$ are in conflict if both of them access a common account and at least one of them performs a write operation, as defined earlier in Equation 6. The read-set and write-set simplify our analysis and improve the accuracy of conflict detection for Solana blocks. We used the beta API of Solana’s mainnet to obtain block details in JSON format with the max-supported transaction version set to 0 [17]. The extracted details are then parsed to obtain the information required for our analysis.
5 Ethereum: An Empirical Study
As shown in Table 1, we selected three distinct historical periods (HPs) based on the major events that impact Ethereum’s network congestion. Each of these periods allow us to assess the blocks in different HPs, giving insights into how major events like popular dApp launches and significant protocol upgrades affect transaction parallelism and conflicts. It also helps us understand the limitations of parallel execution approaches under different network conditions and historical periods.
Table 1: Historical Blocks from Ethereum’s mainnet: 1000 blocks from each historical period.
  CryptoKitties
Deployment (Eck)
    Ethereum 2.0
Merge (Ee2)
    Ethereum Recent
Blocks (Erb)
Block ID of Historical Event 4605167 15537393 21631500
Block Range Before Event 4604664 - 4605166 15536879 - 15537392 21631000 - 21631500
Block Range After Event 4605168 - 4605670 15537394 - 15537907 21631501 - 21632001
Ethereum CryptoKitties Contract Deployment (Eck):
The CryptoKitties [3] game is among the first and the most popular dApps. CryptoKitties was deployed in block 4605167, after which an unexpected spike in transactions caused Ethereum to experience never-before-seen congestion. The workload consists of 500 blocks, each from before and after the deployment of the CryptoKitties smart contract. We can expect this period to receive a high volume of transactions for a contract, consequently leading to congestion at a specific contract, as observed by an earlier study in [13]. Hence, this period is an ideal workload for determining how well the parallel execution approach performs with a large influx of transactions for a contract.
Ethereum 2.0 Merge (Ee2):
Ethereum’s transition from proof-of-work to proof-of-stake consensus took place in block 15537393, called the Ethereum 2.0 merge [5]. This event has changed Ethereum’s consensus mechanism and could have optimized the transaction processing, block validation, and network traffic in general. So in this workload, we try to determine the direct impact of this upgrade on the parallel execution pattern that impacts the transaction throughput and network behavior by comparing blocks before and after the merge.
Ethereum Recent Blocks (Erb):
In addition to the above historical periods, we analyze transactions from the 1000 most recent blocks, ranging from block number 21631000 to 21632001. We selected this range to better understand current transaction access patterns and parallelism under normal conditions when there are no major historical events impacting the network traffic. By examining this workload, we can establish connections between different historical periods tracking Ethereum’s evolution, user access patterns, network congestion over time, and the impact of optimizations on recent blocks. This analysis also helps in developing approaches that take advantage of parallelism efficiently for future network upgrades.
Analysis
As illustrated in Figure 2, Figure 3, and Table 2, the following observations can be made for real-world Ethereum data:
Block SizeETH TransferSmart Contract TxsERC20 TransferIndependent TxsLongest Chain of ConflictsIndependent ETH TxsIndependent SC TxsConflict Families050100150200
CryptoKitties (Eck)
Ethereum 2.0 Merge (Ee2)
Recent Blocks (Erb)
before event - after event before event - after event before event - after event
Block Size 71 - 83 178 - 156 181 - 176
ETH Transfer Txs 35 - 42 66 - 42 65 - 62
Smart Contract (SC) Txs 37 - 41 113 - 114 116 - 114
ERC20 Transfer Txs 12 - 13 11 - 16 43 - 38
Independent Txs (%) 38 (61.87%) - 42 (55.70%) 87 (54.18%) - 55 (39.45%) 92 (51.73%) - 90 (51.82%)
Longest Chain of Conflicts 15 - 15 38 - 42 31 - 27
Ind. ETH Transfer (%) 21 (70.05%) - 24 (63.82%) 32 (62.57%) - 24 (67.72%) 49 (78.29%) - 47 (78.31%)
Independent SC Txs (%) 17 (57.20%) - 19 (52.96%) 56 (56.95%) - 32 (32.52%) 45 (39.82%) - 45 (40.32%)
Conflict Families 30 - 37 56 - 40 72 - 69
71717171353535353737373712121212383838381515151521212121171717173030303083838383424242424141414113131313424242421515151524242424191919193737373717817817817866666666113113113113111111118787878738383838323232325656565656565656156156156156424242421141141141141616161655555555424242422424242432323232404040401811811811816565656511611611611643434343929292923131313149494949454545457272727217617617617662626262114114114114383838389090909027272727474747474545454569696969Transaction CountEck-Before    Eck-AfterEe2-Before    Ee2-AfterErb-Before    Erb-After
Figure 2: Ethereum Historical Periods: analysis based on exclusive access to accounts (addresses) by transactions.
$>$40%$>$50%$>$60%$>$70%$>$80%10100200300400500430430430430348348348348250250250250169169169169939393934124124124122942942942941841841841849797979746464646415415415415316316316316156156156156737373733333333319119119119162626262363636361919191999994764764764763053053053055151515122220047047047047030430430430447474747666600Percentage of Independent Transactions in the BlockBlock CountEck-before   Eck-after   Ee2-before   Ee2-after   Erb-before   Erb-after
Figure 3: Ethereum: the number of blocks where the percentage of independent transactions exceeds the threshold before and after historic event.
Observation-1:
The initial evaluation aims to understand the parallelism by distinguishing between dependent (conflicting) and independent (non-conflicting) transactions, identifying the longest chain of conflicting transactions, and examining conflict families both within and across historical periods.
As shown in Figure 2, transactions per block have increased since $E_{ck}$ HP, with contract transactions rising $\sim$4$\times$ and ETH transfers $\sim$2$\times$. This implies an increased demand for computational resources and an increased adoption of blockchain technology for broad smart contract applications (dApps).
The percentage of independent transactions per block has decreased, particularly post-Ethereum merge, though over 50% remain independent on average. The longest conflict chain comprises $\sim$19–20of the block size, peaking at $\sim$22post-merge and stabilizing at $\sim$16–17in recent blocks. This suggests that even with perfect parallelization of other transactions, including scheduling of conflicting transactions, the maximum speedup is restricted to $\sim$16-17of transactions that must be executed sequentially, the theoretical upper bound on speedup over sequential execution in recent blocks.
Compared to previous HPs, the percentage of independent ETH transfer transactions in recent blocks has increased, whereas the number of independent smart contract transactions has decreased, indicating an upsurge of transactions for specific contracts and diverse user transactions for ETH transfer. However, the rise in conflict families and block sizes from the earlier period to the more recent one suggests that there is a lot of parallelism. This can improve Ethereum’s throughput if parallel transaction execution is employed.
Observation-2:
Calculating the ratio of ETH transfers to smart contract transactions (table in Figure 2), comprising the historical period from Eck to $E_{e2}$ and $E_{rb}$, shows an increased user engagement with contract applications. In Eck HP, the ratio of $\frac{ETH\ transfer}{SC\ Txs}$ is $\frac{38.5}{39}\approx 0.99$, while it is $\sim 0.47$ in $E_{e2}$ and $\sim 0.55$ in $E_{rb}$. This indicates a surge in computational costs over time and the need for parallel transaction execution to improve network throughput.
Observation-3:
To understand how many blocks in each HP have a certain percentage ($>$40%, $>$50%, … $>$80%) of independent transactions and which HP has a higher parallelism compared to others. As shown in Figure 3, we analyzed 1000 blocks of each HP (500 before and 500 after the event).
The number of blocks with more than 40% independent transactions has increased in recent blocks, more than 94% of blocks have at least 40% independent transactions in $E_{rb}$. However, there was a notable decline in independent transactions after each historical event, suggesting a spike in conflicts. Note that more than 50% of the blocks in each HP had more than 50% of independent transactions, while after Ethereum’s merge Ee2 there is a significant drop. The reasons could be an increased congestion for a specific contract (the longest conflict chain increased), a decrease in the number of ETH transfer transactions, and a slight decrease in block size compared to pre-merge, as observed in Figure 3. The number of blocks with a higher percentage of independent transactions could increase if false conflicts are removed using complete read-write access information.
Table 2: Ethereum Recent Blocks (21631001–21631020): analysis based on exclusive access to accounts by transactions.
Block ID
Block
Size
ETH
Transfer
Txs
SC
Txs
ERC20
Transfer Txs
Independent
Txns (%)
Chain of
Conflicts
Independent
ETH Transfer
Txs
Independent
SC Txs
Conflict
Families
Dense
Conflict
Family
21631001 337 130 207 75 146 (43.32%) 56 82 66 116 180
21631002 148 39 109 46 69 (46.62%) 28 34 38 51 89
21631003 82 27 55 35 47 (57.32%) 24 23 27 38 39
21631004 191 65 126 59 92 (48.17%) 50 53 42 73 106
21631005 233 75 158 55 125 (53.65%) 52 69 57 99 127
21631006 154 64 90 34 86 (55.84%) 20 47 40 74 61
21631007 192 67 125 46 101 (52.60%) 28 57 46 75 111
21631008 163 60 103 47 78 (47.85%) 27 46 34 67 86
21631009 177 68 109 55 81 (45.76%) 39 53 31 70 101
21631010 207 86 121 52 81 (39.13%) 52 37 45 68 101
21631011 148 46 102 31 78 (52.70%) 20 41 39 61 87
21631012 134 46 88 32 83 (61.94%) 16 43 40 66 66
21631013 175 51 124 58 93 (53.14%) 34 49 45 72 96
21631014 200 66 134 41 109 (54.50%) 21 54 55 91 85
21631015 138 42 96 40 82 (59.42%) 22 40 42 68 66
21631016 180 68 112 58 76 (42.22%) 34 43 33 61 110
21631017 119 38 81 43 70 (58.82%) 26 34 38 61 50
21631018 230 100 130 46 113 (49.13%) 41 62 52 94 103
21631019 145 47 98 30 84 (57.93%) 19 36 48 67 68
21631020 166 55 111 39 84 (50.60%) 26 45 41 68 87
Average 170 60 110 45 86 (51.70%) 31 46 42 72 91
Observation-4:
Table 2 highlights block-wise trends of the recent historical blocks (the blocks are chosen from the timestamp: January 15, 2025 04:13:23 PM +UTC) where we selected 20 blocks. As shown, most blocks have more than 50% independent transactions, with the highest parallelism in block 21631012, contains 134 transactions out of which 61.94% are independent. The conflict chain is the shortest in this block, with 16 transactions (11.94% of the block size). In particular, most conflicts are from contract transactions; out of 88 contract transactions, 48 (88-40) are conflicting, while only 3 (46-43) are conflicting from ETH transfers. Block 21631010, on the other hand, has the least parallelism, with 210 transactions, only 39.13% of which are independent, and the longest conflict chain involving 25.12% of the block. These blocks have an average of 170 transactions, of which 51.70% are independent. On average, the longest conflict chain takes up 18.23% of the block size.
Table 3: Ethereum recent blocks (21631001–21631020): analysis on additional metrics.
Block ID Block Size SC Deployment Txs Total Conflicts Most Accessed EOA Cumulative EOA Accesses Most Accessed SC Cumulative SC Accesses Chain of Conflicts (%)
Address Access Count Address Access Count
21631001 337 0 191 0x435...d 22 710 0xc02...2 72 576 16.62
21631002 148 0 79 0x435...d 15 318 0xc02...2 29 253 18.92
21631003 82 1 35 0x463...8 8 145 0xdac...7 22 94 29.27
21631004 191 1 99 0x435...d 23 396 0xc02...2 37 343 26.18
21631005 233 1 108 0x435...d 25 554 0xc02...2 50 420 22.32
21631006 154 1 68 0xb5d...1 11 332 0xc02...2 26 236 12.99
21631007 192 1 91 0x435...d 12 396 0xc02...2 47 307 14.58
21631008 163 0 85 0x435...d 20 333 0xdac...7 28 255 16.56
21631009 177 0 96 0x2b3...b 26 338 0xdac...7 40 268 22.03
21631010 207 1 126 0x924...f 34 480 0xdac...7 35 294 25.12
21631011 148 0 70 0x435...d 18 310 0xc02...2 41 265 13.51
21631012 134 1 51 0x435...d 14 303 0xc02...2 29 262 11.94
21631013 175 0 82 0x435...d 13 319 0xc02...2 39 299 19.43
21631014 200 1 91 0x435...d 18 473 0xc02...2 38 384 10.50
21631015 138 0 56 0x435...d 15 314 0xc02...2 29 242 15.94
21631016 180 0 104 0x435...d 26 341 0xdac...7 35 226 18.89
21631017 119 0 49 0x435...d 10 221 0xdac...7 27 137 21.85
21631018 230 2 117 0xeba...1 18 509 0xdac...7 35 334 17.83
21631019 145 1 61 0x435...d 13 303 0xc02...2 26 251 13.10
21631020 166 1 82 0x435...d 13 365 0xc02...2 41 310 15.66
Observation-5:
Table 3 presents additional metrics for block-wise trends of the recent historical blocks 21631001–21631020, as shown in Table 2. These metrics complement the earlier conflict-based analysis in Table 2 by revealing structural dependencies for parallel execution.
From Table 3, we can observe that a small number of EOAs and smart contract accounts are frequently accessed across transactions and become hotspots within each block; e.g., 0x435…d and 0xc02…2 appear as the most accessed addresses in most of the blocks. This indicates a high degree of access skew across different blocks and within each block, which leads to hotspots that can serialize execution in otherwise parallelizable workloads; however, several of these accesses could be just in read mode. For instance, block 21631003 contains only 82 transactions with 145 cumulative EOA accesses, and over 5% of these target a single EOA. This illustrates that even small blocks can exhibit significant access skew. Despite block sizes that vary between 82 and 337 transactions, the average length of the longest conflict chain 18.23% of block size, showing dominant bottlenecks are not due to block size but to contention on shared addresses. Notably, blocks like 21631003 and 21631013 show relatively fewer total transactions; however, they consist of longer dependency chains due to concentrated access to shared smart contracts. These observations suggest that access frequency and locality are stronger indicators of the scope of parallel execution than the raw transaction count.
In Ethereum’s historical blocks, over 50% of the blocks in each HP contained more than 50% independent transactions. The theoretical upper bound on maximum speedup is constrained by the longest conflicting chain, which accounts for approximately 16–17% of block transactions. The change in independent transaction percentages over time and block by block, longest conflict chains, and conflict families indicates that no single parallel execution strategy is optimal for all blocks. Moreover, historical periods show significant shifts in conflict patterns, with smart contract transactions being the primary source of contention. Our observations highlight that we need an adaptive scheduling technique that dynamically chooses the best possible execution strategy and also optimizes overall execution based on real-time block characteristics to maximize throughput and efficiency. Alternatively, a hybrid parallel execution model that leverages conflict information available with transactions with minimum added overhead can maximize the performance of speculative parallel execution and minimize aborts and re-execution overhead.
6 Solana: An Empirical Study
Solana is the first access-specification-aware (read-write aware) blockchain to support parallel execution. Transactions specify the states that are read or written during execution. The Solana Sealevel [21] execution engine executes transactions in parallel using locking-based techniques (read and write locks) to identify independent transactions over multiple iterations [19]. The longest Longest Chain of Conflicts determines the minimum number of iterations required for a block, assuming that a sufficient number of cores are available to fully exploit parallelism. To understand the distribution of conflicts in historical Solana blocks, we analyzed 1000 non-empty blocks from three distinct periods: the old historical period (Sob) from block 61039000 to 61040210, the mid historical period (Smb) from block 205465000 to 205466007, and the recent historical period (Srb) from block 293971000 to 293972009. The Solana block consists of voting and non-voting transactions; we consider non-voting transactions for our analysis.
[…]
  Old Historical
Period (Sob)
  Mid Historical
Period (Smb)
  Recent Historical
Period (Srb)
Block Range 61039000 - 61040210 205465000 - 205466007 293971000 - 293972009
Block Size 519 1972 1249
Non-voting (NV) Txs 252 113 334
Successful NV Txs 215 78 182
Conflicting NV Txs 252 (100%) 101 (87%) 310 (93%)
Independent NV Txs 0 (0%) 12 (13%) 23 (7%)
Chain of Conflicting NV Txs 214 52 119
Conflict Families 3 19 39
Dense Conflict Family Size 233 60 184
Total Conflicts 3664 886 3917
Write-Write Conflicts 3664 676 3751
5195195195192522522522522152152152152522522522520021421421421433332332332332333,66436643{,}6643 , $6643{,}$ 6641,97219721{,}9721 , 97211311311311378787878101101101101121212125252525219191919606060608868868868866766766766761,24912491{,}2491 , 24933433433433418218218218231031031031023232323119119119119393939391841841841843,91739173{,}9173 , 9173,75137513{,}7513 , 751Transaction CountSob  Smb  Srb
Figure 4: Historical Blocks from Solana’s mainnet: 1000 blocks from each historical period and analysis based on read-write sets of non-voting transactions.
Analysis
As illustrated in Figure 4 and Tables 4 and 5, the following observations can be made for real-world Solana blocks:
Observation-1:
As shown in Figure 4, the average block size has increased more than 2$\times$ from the old HP to the recent HP; however, note that this increase is contributed by voting transactions. Non-voting transactions have increased with miner margin, which has seen a deep mid-historical period with increased voting transactions.
Observation-2:
The percentage of successful non-voting transactions has decreased over time, with the success rate of $\sim$85.32in the old historical period to $\sim$69.03in the mid-historical period and to $\sim$54.50in the most recent historical period. This shows that increasing network congestion must have contributed to transaction failures, potentially due to inaccuracies in transaction specifications; the time when specifications are generated by users to the time when executed may differ due to intermediate ongoing execution at the validator nodes. The exact reasons for the increased transaction failures require further analysis, but it could be due to increased network congestion or inaccuracies in the transaction specifications. However, it indicates the limitations and efficiency of read-write aware execution models in high-contention workloads.
Observation-3:
The percentage of independent transactions in Solana blocks is considerably lower than in Ethereum blocks. However, there is a noticeable upward trend, with independent transactions increasing from 0% in the old historical period (Sob) to $\sim$7in the recent historical period (Srb), while the mid-historical period (Smb) recorded $\sim$13%. This suggests a gradual shift toward greater parallelism over time.
Since Solana employs a locking-based multi-iteration parallel execution strategy, the number of conflict families has surged, from just 3 in Sob to 39 in Srb, suggesting that despite high conflicts, multiple independent subsets of transactions can still be executed in parallel. Each subset was executed in parallel with the others, enhancing the execution efficiency.
The longest chain of conflicting transactions, relative to the total number of non-voting transactions in a block, has seen a substantial decline. Specifically, the longest conflict chain has been reduced by 2.3$\sim$ 2.3$\times$ 2.3 ×. The longest chain of conflicting transactions decreased from 84.92% in Sob HP to 46.01% in Smb HP and further to 35.62% in the most recent Srb HP. The number of transactions within the most densely conflicted family has also seen downward trends. It suggests more distributed conflicts and the possibility of improved parallel execution with more granular bottlenecks in transaction execution. Showing that transaction access patterns have changed over time, consequently improved throughput of Solana’s read-write aware execution model.
Observation-4:
Note that the majority of conflicts are from write sets in historical blocks, accounting for $\sim$100in the old historical period, which decreased by $\sim$4.24($\sim$95.76%) in recent blocks. This suggests that any approach that could minimize write-write conflicts could significantly enhance Solana’s throughput. A potential solution could be to adopt a multi-version data structure, similar to the one employed in Block-STM [7], which allows parallel execution while minimizing write-write contention.
Table 4: Solana recent historical blocks (314184230–314184249): analysis based on read-write sets of transactions.
Block ID
Block
Size
Non-voting
(NV) Txs
Successful
NV-Txs
Independent
NV-Txs (%)
Chain of
Conflicts
Conflict
Families
Dense Conflict
Family Size
Total
Conflicts
W-W
Conflicts
314184230 2281 590 454 20 (3.39%) 386 32 527 16612 16145
314184231 1537 518 170 22 (4.25%) 326 33 374 17116 17071
314184232 2090 515 477 34 (6.6%) 326 46 407 8955 8871
314184233 1689 335 298 2 (0.6%) 238 6 301 5903 5566
314184234 1844 431 276 3 (0.7%) 226 6 419 8619 5762
314184235 1304 72 40 6 (8.33%) 23 13 29 212 212
314184236 2075 446 403 41 (9.19%) 239 59 283 4166 3803
314184237 1722 330 304 11 (3.33%) 232 20 283 3184 3184
314184238 1439 586 372 30 (5.12%) 375 42 519 4890 4739
314184239 2026 517 377 17 (3.29%) 366 28 455 10900 5906
314184240 2331 473 420 41 (8.67%) 266 54 355 6738 4073
314184241 1863 498 262 16 (3.21%) 304 22 460 5614 4715
314184242 1723 327 188 16 (4.89%) 197 23 295 2715 2102
314184243 1579 430 192 6 (1.4%) 285 9 417 7672 5887
314184244 2338 644 342 28 (4.35%) 299 38 533 12773 7727
314184245 1568 544 348 19 (3.49%) 325 28 490 5018 4332
314184246 2239 485 312 13 (2.68%) 314 21 436 11463 4036
314184247 1484 477 233 16 (3.35%) 309 19 450 8480 4009
314184248 2620 781 308 26 (3.33%) 320 38 394 15345 11245
314184249 1711 535 210 11 (2.06%) 322 18 476 9763 7219
Average 1873 477 299 19 (4%) 284 28 395 8307 6330
Observation-5:
Table 4 highlights block-by-block analysis in recent blocks (the blocks are chosen from the same period as Ethereum: January 15, 2025 04:13:23 PM +UTC). As shown, the size of the block varies with significant margin, while the number of non-voting transactions is in the range from $\sim$72 to $\sim$781 per block, indicating increased voting activity in the network with more participating validator nodes over past historical periods. However, the independent transaction percentage varies from a minimum of $\sim$0.6in block 314184233 to a maximum of $\sim$9.19in block 314184236, with an average of $\sim$4%, which is considerably lower compared to Ethereum blocks. Additionally, in all the aforementioned blocks, the majority of conflicts originate from write sets. The average longest conflict chain consists of 284 transactions, accounting for $\sim$59of the non-voting transactions in the block, further emphasizing the possibility of write-write conflict optimization in pursuit of efficient parallel execution.
Table 5: Solana recent historical blocks (314184230–314184249): additional analysis of read-write sets and account access patterns.
Block ID Non-Voting (NV) Txs Largest Rset Size Average Rset Size Largest Wset Size Average Wset Size Accounts Accessed Most Read Account Most Written Account
Address Count Address Count
314184230 590 21 6 34 8 2573 Comp...1 449 GZ...ib 155
314184231 518 21 8 27 10 1172 Comp...1 485 J3...S7 110
314184232 515 26 6 42 6 2309 Comp...1 356 GZ...ib 95
314184233 335 18 6 19 7 1676 Comp...1 255 GZ...ib 90
314184234 431 22 8 34 9 2420 Comp...1 403 GZ...ib 75
314184235 72 20 8 24 9 437 Comp...1 67 Ce...iM 16
314184236 446 21 6 37 7 2326 $1$...1 323 Ce...iM 40
314184237 330 22 7 35 9 1861 $1$...1 289 Ce...iM 40
314184238 586 21 8 35 10 2816 Comp...1 451 oQ...d7 44
314184239 517 27 8 35 9 2623 Comp...1 403 Ce...iM 77
314184240 473 38 7 38 8 2649 $1$...1 368 FR...Qs 44
314184241 498 24 9 35 11 2283 Comp...1 424 7e...Aw 46
314184242 327 26 9 43 14 2814 Comp...1 263 HW...oY 29
314184243 430 21 9 50 11 2191 Comp...1 380 4Y...QU 61
314184244 644 24 8 42 9 2906 Comp...1 565 Ce...iM 74
314184245 544 24 8 46 11 2933 Comp...1 419 Ce...iM 33
314184246 485 24 8 46 10 2629 Comp...1 394 VN...nP 40
314184247 477 24 9 50 12 2779 Comp...1 401 VN...nP 35
314184248 781 22 7 46 7 2730 Comp...1 722 6r...hE 44
314184249 535 38 8 46 12 2492 Comp...1 510 9D...QX 84
Observation-5:
Table 5 presents additional metrics for block-wise trends of recent Solana blocks 314184230–314184249, analyzed through their transaction-level read-write sets. These metrics complement earlier contention or scheduling analyses by characterizing structural features that influence Solana’s parallel execution model.
From Table 5, we observe consistently high reuse of specific accounts across blocks, especially the ComputeBudget1111... account, which is read in 17 out of the 20 blocks and dominates access frequency (e.g., 722 reads in block 314184248). Similarly, certain smart contract accounts like CebN5W... and VN1RtW... recur frequently as the most written accounts. This illustrates a access skewness, where a few accounts become systemic hotspots across blocks. Interestingly, the number of non-voting transactions varies significantly, from 72 (block 314184235) to 781 (block 314184248), however access contention remains prominent regardless of block size. For example, both blocks 314184240 and 314184249 show the largest Rset and Wset sizes (up to 38 and 46 accounts, respectively), reflecting extensive data dependencies. These structural patterns suggest that Solana’s optimistic parallel execution must still grapple with dependency bottlenecks arising from frequent access to shared accounts. Overall, the results indicate that access locality and skew, particularly with respect to protocol-level accounts (e.g., ComputeBudget1111…), for system-level reads or writes play a dominant role in shaping the effective parallelism, often more than block size or average set size alone.
With these observations, we conclude that independent transactions remain very low on the Solana network in all three historical periods, with an average of $\sim$4in recent blocks, while write-write conflicts dominate and contribute to $\sim$76.2of all conflicts within a given block. The longest chain of conflicting transactions has decreased significantly, from 84.92% to 35.62%, from the old historical period to the recent historical period, but the number of conflict families has increased from 3 to 39, indicating more granular bottlenecks (conflicts) and increased parallelism. Furthermore, the success rate of non-voting transactions has dropped from $\sim$85.32to 54.50%, which highlights the need for adaptive or hybrid-execution strategies that exploit access specifications efficiently to improve throughput and reduce failure rates.
7 Conflicts on Ethereum versus Solana
As shown in Figure 5, Solana’s larger block size (352, non-voting transactions) compared to Ethereum (170) highlights a fundamental architectural difference between the two blockchains. Solana supports high throughput and parallel execution, leading to higher raw data per block due to its read-write aware execution. Ethereum, in contrast, structures its blocks around finalized, gas-accounted transactions, reflecting a design optimized for decentralization and resource efficiency.
This distinction extends to conflict behavior: Ethereum exhibits a higher percentage of independent transactions (86, 50.59%) compared to Solana (29, 8.24%), reflecting a more modular transaction structure with a greater focus on isolated transaction processing. Ethereum also has a higher percentage of conflict families (72, 42.35%) compared to Solana (48, 13.64%), further supporting the idea that its transaction structure tends to experience more isolated contention points. In contrast, Solana’s design, optimized for high throughput, results in longer chains of conflicts (109, 30.97%) compared to Ethereum (31, 18.24%), indicating deeper interdependencies between transactions. Although Solana has fewer conflict families, its dense conflict family size (176, 50%) is closer to that of Ethereum (91, 53.53%), suggesting a similar level of concentration in transactional contention within specific resource sets across both networks. This observation highlights a limit on the maximum achievable parallel efficiency in both blockchains, indicating that while both platforms are optimized for high throughput, they face inherent constraints related to transaction dependencies and resource contention. These patterns, which span block composition, transaction independence, and conflict complexity, illustrate how Ethereum and Solana make different trade-offs in their pursuit of efficient execution.
Block SizeIndependent TxnsLongest Chain of ConflictsConflict FamiliesDense Conflict Family Size050100150200250300350400170170170170868686863131313172727272919191913523523523522929292910910910910948484848176176176176Transaction CountEthereumSolana
Figure 5: Ethereum versus Solana in recent historical period.
In summary, it is crucial to understand the conflict matrices for transactions on blockchains for efficient parallel execution. Both Ethereum and Solana exhibit significant parallel execution potential; they differ in key aspects of conflicts, available parallelism, and potential blockers for efficient execution. Ethereum, with lower conflict rates with more independent transactions that could result in higher parallelism. Showing potential for execution efficiency and lower abort rates in optimistic execution. Solana, on the other hand, exhibits higher conflict rates, particularly due to write-write conflicts, resulting in more granular congestion or conflicts, highlighting the limitations of its current execution model. However, both blockchains could be further optimizations when it comes to execution.
Despite Solana’s higher transaction throughput compared to Ethereum on its mainnet, it faces the challenge of increasingly common transaction failures. Likewise, since Ethereum still executes transactions sequentially; there is ongoing research in parallel execution approaches for EVM [8, 12, 14, 15, 18] inspired by software transactional memory that could handle contention more effectively for Ethereum. Given current trends, we believe that both networks (Ethereum and Solana, including other popular EVM and SVM-based chains) would benefit from adaptive and hybrid scheduling techniques to exploit parallelism in favor of higher throughput. Solana, in particular, requires innovations to mitigate write-write conflicts, potentially through the adoption of multi-version data structures.
8 Concluding Remarks
Blockchains typically process transactions in a strict order, one after another, to ensure that all nodes reach identical states. This approach guarantees consistency, but limits throughput. Identifying which transactions truly depend on each other, i.e., conflict, can help maximize blockchain performance. The EVM processes transactions sequentially, without advance knowledge of what state each transaction will access. The SVM, on the other hand, requires clients to specify upfront what states (accounts) a transaction will read or write.
Our analysis of Ethereum blocks revealed something interesting: consistently across all time periods, over 50% of the transactions are completely independent and could theoretically be executed in parallel. In recent blocks, 51.7% of transactions are independent (can be parallelized). The longest conflict chain, those that must run sequentially, is only about 16-17% of the block size. Ethereum smart contract transactions in recent blocks, show higher conflict rates ($\frac{110-42}{110}\approx 0.61$) than simple ETH transfers ($\frac{60-46}{60}\approx 0.23$).
In contrast, Solana blocks exhibit distinct conflict patterns. From the old to the recent historical period, the longest conflict chain shortened (84.92% to 35.62%), while conflict families increased (3 to 39), indicating a shift toward fine-grained parallelism. However, recent blocks show consistently low independence (avg. $\sim$4%), with write-write conflicts comprising most contention ($\sim$76%). Additionally, the success rate of non-voting transactions dropped from $\frac{215}{252}\approx 0.85$ to $\frac{299}{477}\approx 0.62$, highlighting the limitations of Solana’s current read-write aware scheduling under high contention workloads.
All of this suggests that block-by-block variation is significant, necessitating adaptive execution strategies. Moreover, this analysis in the paper confirms what we suspected: traditional sequential execution is leaving enormous performance potential untapped, which can be exploited with proper conflict detection and parallel execution techniques in both the Ethereum and Solana Virtual Machines.